Cicchetti Corporation uses customers served as its measure of activity. The following report compares the planning budget to the actual operating results for the month of December:

Cicchetti Corporation
Comparison of Actual Results to Planning Budget
For the Month Ended December 31
Actual
Results
Planning Budget Variances
  Customers served 31,000   30,000  
  Revenue (3.50q) $ 108,800 $ 105,000 $ 3,800 F
  Expenses:
     Wages and salaries ($23,000 + $1.20q) 60,200 59,000 1,200 U
     Supplies ($0.60q) 16,100 18,000 1,900 F
     Insurance ($4,900) 4,900 4,900 0
     Miscellaneous expense ($3,900 + $.40q) 14,000 15,900 1,900 F
     Total expense 95,200 97,800 2,600 F
  Net operating income $ 13,600 $ 7,200 $ 6,400 F
Required:
1.

Prepare a report showing the company’s revenue and spending variances for December. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)
Answer #1
Revenue and spending Variance
Actual Results Revenue and spending Variance Flexible Budget
Customer served 31000 31000
Revenue 108800 300 F 108500
Expenses
   Wages and salaries 60200 None 60200
   Supplies 16100 2500 F 18600
   Insurance 4900 None 4900
   Miscellaneous Expenses 14000 2300 F 16300
Total expenses 95200 4800 F 100000
Net Operating Income 13600 5100 F 8500
Flexible Budget
Revenue = 3.50 * 31,000 = $ 108,500
Wages and salaries = $ 23,000 + ( 1.20 * 31,000 ) = $ 60,200
Supplies = 0.60 * 31000 = $ 18,600
Miscellaneous expenses = $ 3,900 + ( .40 * 31,000 ) = $ 16,300